OpenStudy (anonymous):

1/10(3x-7) = 3/10x + 5 just need to check my work

OpenStudy (amistre64):

is the the final result that your checking?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got no solution

OpenStudy (anonymous):

I'm assuming the problem is: \[\frac{1}{10(3x-7)}=\frac{3}{10x+5}\] Cross multiply: \[10x+5=30(3x-7)\] \[10x+5=90x-210\] \[80x=215\] \[x=\frac{215}{80}\] \[x=\frac{43}{16}\]

OpenStudy (amistre64):

well, lets step thru it, first, is this equational stuff? \[\frac{1}{10(3x-7)} = \frac{3}{10x} + 5\]

OpenStudy (anonymous):

no i really need to learn to put it down the way you do the 1is over the 10

OpenStudy (anonymous):

and (3x-7)

OpenStudy (amistre64):

like this? \[\frac{1}{10}(3x-7) = \frac{3}{10x} + 5\]

OpenStudy (anonymous):

The method used is \frac{numerator}{denominator}.

OpenStudy (amistre64):

the equation editor button at the bottom can be useful, but its might be difficult to use

OpenStudy (anonymous):

yes that is right

OpenStudy (amistre64):

is that x on the right side under or beside the 3/10? \[\frac{1}{10}(3x-7) = \frac{3}{10}x + 5\]

OpenStudy (anonymous):

that is right it is beside the \[3/10\]

OpenStudy (amistre64):

ok \(\cfrac{1}{10}(3x-7) = \cfrac{3}{10}x + 5\) ; lets multiply it all by 10, do you see why? \(\cfrac{1(10)}{10}(3x-7) = \cfrac{3(10)}{10}x + 5(10)\) \(3x-7 = 3x + 50\)

OpenStudy (amistre64):

and yes, when we have the same value of x on both sides it dissapears and we get no solution

OpenStudy (amistre64):

say 3x = 4 for example... 4-7 = 4+50 ?? no

OpenStudy (amistre64):

the only time it would give a solution is when you have the same thing on both sides; then the solution is "all numbers"

OpenStudy (anonymous):

right i see

OpenStudy (anonymous):

the reason we used ten is that it is LCD and then we divided it by the LCD and do we do that every time?

OpenStudy (amistre64):

dont worry about LCM, i never do. see what you need to do to clear the fractions; by that i mean turn them into "1" 3/5 would be cleared if that 5 on the bottom was gone right? 3(5)/5 = 3 * 5/5 = 3*1 = 3 to clear the fractions, do them one at a time if there are different denominators

OpenStudy (amistre64):

3/4 + 7/3 = x ; clear the fractions 3(4)/4 + 7(4)/3 = (4)x 3 + 28/3 = 4x ; and again 3(3) + 28(3)/3 = 4(3)x 9 + 28 = 12x 37 = 12x ; divide off the 12 37/12 = x reduce as wanted